# Diffusion Reward: Learning Rewards via Conditional Video Diffusion

### v2

@@ 5.5 Ablation Studies
Conditional entropy with diffusion model. As our method has demonstrated better performance than VIPER, we take a further step to see the joint effect of the reward types and the video prediction models. Specifically, we systematically evaluate all possible combinations of conditional entropy and log-likelihood as reward signals, each paired with either diffusion-based or transformer-based video prediction models. Note that the same vector-quantized encoder is used for all models, ensuring the variations are solely attributable to the chosen video prediction model and reward. The outcomes consistently align with the observations delineated in Figure 2, indicating two-fold conclusions: (1) video diffusion models are more adept at capturing the complex distribution of expert videos in complex tasks, thus resulting in more informative rewards; (2) employing conditional entropy as rewards prove more productive in RL exploration than using log-likelihood, partially owing to the more generalizable reward inference on trajectories unseen in reward pretraining stage.
@@
Denoising steps. The number of timesteps involved in the reverse process governs the quality and diversity of generated frames [33]. This study seeks to investigate its impact on the derived reward and subsequent RL performance by gradually increasing the number of denoising steps from 2 to 20. The findings indicate that an intermediate choice, approximately around 10 steps, achieves the best performance. This suggests that an intermediate choice balances generative quality and diversity well, thereby producing effective rewards for RL exploration. Furthermore, we observe that the speed of reward inference declines with an increase in denoising steps. For instance, the Frames Per Second (FPS) during RL training drops from approximately 100 with 2 steps to 60 with 20 steps in Adroit tasks with NVIDIA A40 , suggesting the importance of adopting advanced techniques to expedite the diffusion process.
@@
Sampling noise in diffusion process. We hypothesize that the introduction of randomness in the diffusion process holds the potential to enhance RL exploration, akin to the stochastic characteristics inherent in maximum entropy RL [46]. To substantiate this point, we design a variant of Diffusion Reward wherein the sampling noise is deliberately set as 0 during the reverse process, ensuring the reward becomes deterministic given the identical historical observations. The outcomes show a discernible degradation in performance when employing a deterministic reward. Notably, this decline in performance aligns with results observed when combining the diffusion model with log-likelihood rewards, wherein the rewards are also deterministic. Consequently, our findings demonstrate that the inherent randomness of Diffusion Reward from the reverse process indeed contributes to RL exploration.
@@
Reward coefficient $\alpha$. The reward coefficient $\alpha$ determines the relative importance of the conditional entropy reward against novelty-seeking. We investigate the effect of this parameter by gradually decreasing the value of $\alpha$ from 1 to 0 and repeatedly train the RL agent with the remaining settings identical. The results show that $\alpha$ around 0.95 achieves the best performance, while too-large ones (akin to RND only) and too-small ones (akin to no RND) exhibit significant performance drops. This suggests the domination of Diffusion Reward may still result in getting stuck to local optima, while our proposed reward effectively helps RL agent to narrow down the wide intended exploration space of novelty-seeking rewards.
@@
Context length. The number of historical frames determines the extent of temporal information being encoded during video diffusion, thus influencing the generating process of video diffusion and induced reward inference. To investigate its effect on downstream RL, we test different choices of context length. The results suggest that opting for 1 or 2 historical frames proves sufficient to generate highly effective rewards, owing to the robust generative capabilities inherent in the diffusion model. Interestingly, a marginal decline in performance is observed when the context length is extended to 4 or 8 frames. This phenomenon may be attributed to potential overfitting to expert trajectories, resulting in inferred rewards that exhibit suboptimal generalization to previously unseen trajectories.
@@ 6 Conclusion
In this work, we propose Diffusion Reward, a novel framework that extracts dense rewards from a pre-trained conditional video diffusion model for reinforcement learning tasks. We first pre-train a video diffusion model on expert videos and find that the entropy of the predicted distribution well discriminates the expert-level trajectories and under-expert-level ones. Therefore, we use its standardized entropy, plus the exploration reward and the sparse environmental reward, as an informative reward signal. We evaluate Diffusion Reward over 10 visual robotic manipulation tasks from MetaWorld and Adroit and observe prominent performance improvements over 2 domains. We also demonstrate that our pre-trained reward could directly produce reasonable rewards in unseen tasks, largely surpassing baseline methods. This underscores the potential of large-scale pretrained diffusion models in reward generalization.
@@
Future work will leverage larger diffusion models sourced from a wider dataset to solve diverse simulation and real-world tasks. The incorporation of additional modalities, such as language, will be explored to augment the generalization capabilities of Diffusion Reward. In addition, enhancing the diffusion-based reward itself, including strategies to balance entropy reward and exploration reward and the estimations of conditional entropy, holds promise for yielding better outcomes.
Appendix A Implementation Details
@@
In this section, we provide further implementation details on Diffusion Reward and baselines. Note that all methods use the same RL backbone and maintain all settings except reward pretraining (if exist) identical.
A.1 Diffusion Reward Implemenatation
@@
Codebase. Our codebase of VQ-GAN [10] is built upon the implementation in https://github.com/dome272/VQGAN-pytorch, which provides clean code structure and show fast inference speed. The codebase of VQ-Diffusion [12] is built upon the official implementation, which is publicly available on https://github.com/microsoft/VQ-Diffusion. For the downstream RL, we adopt the official implementation of DrQv2 [39] as RL backbone, which is publicly available on https://github.com/facebookresearch/drqv2, and the implementation of RND as exploration reward available on https://github.com/jcwleo/random-network-distillation-pytorch.
@@
Network architectures. The major network architectures employed in Diffusion Reward follow the original implementation provided by the codebase above (refer to corresponding papers for more details), except for modifications performed in the conditional video diffusion part. Specifically, each historical frame, encoded by the encoder $E$ and quantizer $Q$ learned with VQ-GAN, is tokenized ($8\times 8$) by the condition network, concatenated with others ($2\times 8\times 8$), fed into embedding networks with a dimension of 1024. The resulting condition embedding with a dimension of $128\times 1024$ is passed to subsequent conditional diffusion.
@@
Hyperparameters. We list the important hyperparameters of VQ-GAN, VQ-Diffusion, DrQv2 with Diffusion Reward in Table 45, and 6, respectively.
@@
Entropy estimation details. As described in Section 4.2, the variational bound of conditional entropy in Eq. (1) can be estimate by Eq. (3). Such estimation is realized with the closed-form distribution [32, 20] (e.g., discrete multivariate distribution) in this work. Specifically, the variational bound of conditional log-likelihood in Eq. (1) can be simplified following [32], resulting in our estimation of entropy reward $r^{\mathrm{ce}}$ as follows:
@@
where $D_{KL}$ denotes the Kullback–Leibler divergence, $p(\tilde{\bm{z}}_{k}^{T})$ follows the prior distribution of random noise at timestep $T$, and $\tilde{\bm{z}}_{k}^{0:T}\sim p_{\theta}(\bm{z}_{k}^{0:T}|\bm{z}_{c})$ represents the denoised samples via inverse diffusion process. We set $M=1$ to ensure high reward inference speed while retaining its discrimination on expert-like and -unlike behaviors. We present more analyses in Section D.
A.2 Baselines Implementations
RND implementation. This baseline combines the sparse environmental reward with the RND exploration reward, which is equivalent to setting the reward coefficient $\alpha$ as 1 in Diffusion Reward. To this end, we implement this baseline by simply removing the entropy reward and keeping other settings identical.
VIPER implementation. We implement their adopted VideoGPT based on the official code provided in https://github.com/wilson1yan/VideoGPT with clean GPT implementation from https://github.com/karpathy/minGPT. The calculation of video prediction (log-likelihood) rewards follows the official JAX implementation provided in https://github.com/Alescontrela/viper_rl, which uses ’teacher-forcing’ practice [2] (where ground truth context is provided for each step) for fast inference speed. The coefficient between video prediction reward and RND exploration reward is set as 0.5 following their paper [9].
AMP implementation. The implementation is based on the official code in https://github.com/xbpeng/DeepMimic and re-implementation in https://github.com/med-air/DEX. The encoder consists of three 32-channel convolutional layers interpolated with ReLU activation. The discriminator is implemented as a 3-layer MLP with hidden dimensions of 256 and Tanh activation.
Appendix B Task Descriptions
We select 7 gripper manipulation tasks from MetaWorld [40] and 3 dexterous manipulation tasks from Adroit [28], as visualized in Figure 10. The tasks are widely used in visual RL [13] and are chosen to be diverse in objects and manipulating skills. All tasks render $64\times 64$-dimensional RGB as the agent’s observation and produce sparse environmental rewards. According to task complexity, we collect 20 expert videos for each MetaWorld task and 50 for each Adroit task. We describe each task below:
@@ Appendix B Task Descriptions
Figure 11: Ablations on diffusion process. The results in (a) suggest that overmuch denoised samples (i.e., 4) may hinder the exploration due to the low variance of estimated entropy. This is further verified in (b) where an appropriate choice of sampling noise scale results in more productive explorations. Results are means of 3 seeds with standard errors (shaded area). Red is our default.
Appendix C More Video Prediction Results
Qualitative results. We present the comparison between expert videos and prediction results in Figure 14. We find that the adopted video diffusion model is able to capture the complex distribution of expert videos from pretraining data and generalize well to unseen expert videos. Interestingly, we also observe that the colored target points in Door Close and Reach are sometimes mispredicted, which may explain the relatively slow exploration at the initial RL training stage, suggesting that a more powerful video diffusion model could be used to further improve the reward quality.
Quantitative analysis. We compare the video prediction quality between Diffusion Reward and VIPER in terms of three video metrics, SSIM, PSNR, and LPIPS. Results are shown in Table 1, verifying that Diffusion models hold a stronger generalization ability on unseen real robot and simulation trajectories than VideoGPT and thus produce more informative rewards. For real robot videos, we use $16\times 16$ codes to represent images due to the scene complexity.
@@ Appendix C More Video Prediction Results
Appendix D More Ablations
We conduct more ablations on our proposed Diffusion Reward in this section. Results are aggregated over Door and Hammer from Adroit with 3 random seeds.
Sparse environmental reward. The sparse environmental reward $r^{\mathrm{spar}}$ is integrated into Eq. (5), as the environmental supervision of completion is helpful for RL. We remove $r^{\mathrm{spar}}$ to study its effect. The results in Figure 12 show that the performances decrease dramatically without sparse environmental rewards, indicating that the signal of task completion is necessary for solving complex manipulation tasks. Interestingly, more favorable performance is observed in the Door task. We attribute this to the overlap supervision of RND reward and sparse reward, i.e., exploring novel states (door opening) is partially equivalent to completing the task.
Diffusion process. Recall that, in Eq. (3), we perform inverse process for $M$ times, and use the generated $M$ samples to estimate the conditional entropy as rewards. The results in Figure 11(a) show that the effect of the number of denoised samples has a slight influence on RL performance when increasing from 1 to 3. However, the learning progress gets stuck in the middle state with 4 denoised samples, though the asymptotic performance is still satisfactory. We posit that this is due to the low variance of estimated entropy, which may lead to more weight on exploitation instead of exploration.
@@ Appendix D More Ablations
Meanwhile, we present the time efficiency of different numbers of diffusion processes with NVIDIA A40 in Table 3, where the Frames Per Second (FPS) decreases from 87.7 to 45.9 when the number of denoise samples increases from 1 to 4. It suggests that using 1 denoised sample is sufficient to provide informative rewards and retain high inference speed.
@@
Appendix E More Generalization Experiments

### v3

@@ 5.5 Ablation Studies
As our method has demonstrated better performance than VIPER, we take a further step to see the joint effect of the reward types and the video prediction models. Specifically, we systematically evaluate all possible combinations of conditional entropy and log-likelihood as reward signals, each paired with either diffusion-based or transformer-based video prediction models. Note that the same vector-quantized encoder is used for all models, ensuring the variations are solely attributable to the chosen video prediction model and reward. The outcomes consistently align with the observations delineated in Figure 3, indicating two-fold conclusions: (1) video diffusion models are more adept at capturing the complex distribution of expert videos in complex tasks, thus resulting in more informative rewards; (2) employing conditional entropy as rewards prove more productive in RL exploration than using log-likelihood, partially owing to the more generalizable reward inference on trajectories unseen in reward pretraining stage.
Figure 11: Ablations. Success rate curves for ablated versions of Diffusion Reward, aggregated over Door and Hammer from Adroit. (a) We test the combinations of generative models and rewards to show the benefit of estimating conditional entropy with the diffusion model. (b) We ablate the choice of the number of denoising steps. (c) We demonstrate that the inherent randomness of Diffusion Reward from the reverse process helps RL exploration. (d) We ablate the choice of reward coefficient. (e) We test the effect of the number of conditional frames. Results are means of 3 seeds with the standard error (shaded area). Red is our default.
Denoising steps.
@@
The number of timesteps involved in the reverse process governs the quality and diversity of generated frames [43]. This study seeks to investigate its impact on the derived reward and subsequent RL performance by gradually increasing the number of denoising steps from 2 to 20. The findings indicate that an intermediate choice, approximately around 10 steps, achieves the best performance. This suggests that an intermediate choice balances generative quality and diversity well, thereby producing effective rewards for RL exploration. Furthermore, we observe that the speed of reward inference declines with an increase in denoising steps, suggesting the benefit of adopting advanced techniques to expedite the diffusion process if requiring more denoising steps.
Sampling noise in diffusion process.
@@
We hypothesize that the randomness in the diffusion process can accelerate RL exploration, akin to the stochastic characteristics of maximum entropy RL [60]. To substantiate this point, we design a variant of Diffusion Reward wherein the sampling noise is deliberately set as 0 during the reverse process, ensuring the reward becomes deterministic given the identical historical observations. The outcomes show a discernible degradation in performance when employing a deterministic reward. Notably, this decline in performance aligns with results observed when combining the diffusion model with log-likelihood rewards, wherein the rewards are also deterministic. Consequently, our findings demonstrate that the inherent randomness of Diffusion Reward from the reverse process indeed contributes to RL exploration.
Reward coefficient $\alpha$.
@@
The reward coefficient $\alpha$ determines the relative importance of the conditional entropy reward against novelty-seeking. We investigate the effect of this parameter by gradually decreasing the value of $\alpha$ from 1 to 0. The results show that $\alpha$ around 0.95 achieves the best performance, while too-large ones (akin to RND only) and too-small ones (akin to no RND) exhibit significant performance drops. This suggests the domination of Diffusion Reward may still result in getting stuck to local optima, while our proposed reward effectively helps the RL agent to narrow down the wide exploration space of novelty-seeking rewards.
Context length.
@@
The number of historical frames determines the extent of temporal information being encoded during video diffusion, thus influencing the generating process of video diffusion and induced reward inference. To investigate its effect on downstream RL, we test different choices of context length. The results suggest that opting for 1 or 2 historical frames proves sufficient to generate highly effective rewards, owing to the robust generative capabilities inherent in the diffusion model. Interestingly, a marginal decline in performance is observed when the context length is extended to 4 or 8 frames. This phenomenon may be attributed to potential overfitting to expert trajectories, resulting in inferred rewards that exhibit suboptimal generalization to previously unseen trajectories.
@@ 6 Conclusion
In this work, we propose Diffusion Reward, a novel framework that extracts dense rewards from a pretrained conditional video diffusion model for RL tasks. We first pretrain a video diffusion model using expert videos and observe that the entropy of the predicted distribution well discriminates the expert-level trajectories and under-expert-level ones. Consequently, we use its standardized entropy, plus the exploration reward and the sparse environmental reward, as an informative reward signal. We evaluate Diffusion Reward across 10 visual robotic manipulation tasks from MetaWorld and Adroit and observe prominent performance improvements over 2 domains. Additionally, we present results of reasonable rewards generated in real robot tasks and associated promising offline RL performance, highlighting the real-world applicability of our method. Furthermore, we demonstrate that our pretrained rewards can instruct RL to solve unseen tasks successfully and effectively, largely surpassing baseline methods. This underscores the potential of large-scale pretrained diffusion models in reward generalization.
Limitation and future work.
@@
Future work will leverage larger diffusion models sourced from a wider dataset to solve diverse simulation and real-world tasks. The incorporation of additional modalities, such as language, will be explored to augment the generalization capabilities of Diffusion Reward. In addition, enhancing the diffusion-based reward itself, including strategies to balance entropy reward and exploration reward, the estimations of conditional entropy, and accelerating reward inference speed, holds promise for yielding better outcomes.
Acknowledgements
This work is supported by the National Key R&D Program of China (2022ZD016 1700). We thank all the reviewers for their insightful comments.
Appendix 0.A Implementation Details
@@
In this section, we provide further implementation details on Diffusion Reward and baselines. Note that all methods use the same RL backbone and sparse task reward. We maintain all settings except reward pretraining (if exist) identical.
0.A.1 Diffusion Reward Implemenatation
Codebase.
@@ 0.A.1 Diffusion Reward Implemenatation
Our codebase of VQ-GAN [12] is built upon the implementation in https://github.com/dome272/VQGAN-pytorch, which provides clean code structure and shows fast inference speed. The codebase of VQ-Diffusion [13] is built upon the official implementation, which is publicly available on https://github.com/microsoft/VQ-Diffusion. For the downstream RL, we adopt the official implementation of DrQv2 [51] as RL backbone, which is publicly available on https://github.com/facebookresearch/drqv2, and the implementation of RND as exploration reward available on https://github.com/jcwleo/random-network-distillation-pytorch.
Network architectures.
@@
The major network architectures employed in Diffusion Reward follow the original implementation provided by the codebase above (refer to corresponding papers for more details), except for modifications performed in the conditional video diffusion part. Specifically, each historical frame, encoded by the encoder $E$ and quantizer $Q$ learned with VQ-GAN, is tokenized ($8\times 8$) by the condition network, concatenated with others ($2\times 8\times 8$), fed into embedding networks with a dimension of 1024. The resulting condition embedding with a dimension of $128\times 1024$ is passed to subsequent conditional diffusion.
Hyperparameters.
@@
We list the important hyperparameters of VQ-GAN, VQ-Diffusion, DrQv2 with Diffusion Reward in Table 45, and 6, respectively.
Entropy estimation details.
@@
As described in Section 4.2, the variational bound of conditional entropy in Eq. (1) can be estimate by Eq. (3). Such estimation is realized with the closed-form distribution [42, 20] (e.g., discrete multivariate distribution) in this work. Specifically, the variational bound of conditional log-likelihood in Eq. (1) can be simplified following [42], resulting in our estimation of entropy reward $r^{\mathrm{ce}}$ as follows:
@@
where $D_{KL}$ denotes the Kullback–Leibler divergence, $p(\tilde{\bm{z}}_{k}^{T})$ follows the prior distribution of random noise at timestep $T$, and $\tilde{\bm{z}}_{k}^{0:T}\sim p_{\theta}(\bm{z}_{k}^{0:T}|\bm{z}_{c})$ represents the denoised samples via inverse diffusion process. We set $M=1$ to ensure high reward inference speed while retaining its discrimination on expert-like and -unlike behaviors. We present more analyses in Section 0.D.1.
Figure 12: Task descriptions. (left) 10 seen training tasks from MetaWorld and Adroit. (right) 8 unseen tasks from MetaWorld.
0.A.2 Baselines Implementations
RND implementation.
This baseline combines the sparse environmental reward with the RND exploration reward, which is equivalent to setting the reward coefficient $\alpha$ as 1 in Diffusion Reward. To this end, we implement this baseline by simply removing the entropy reward and keeping other settings identical.
VIPER implementation.
We implement their adopted VideoGPT based on the official code provided in https://github.com/wilson1yan/VideoGPT with clean GPT implementation from https://github.com/karpathy/minGPT. The calculation of video prediction (log-likelihood) rewards follows the official JAX implementation provided in https://github.com/Alescontrela/viper_rl, which uses ’teacher-forcing’ practice (where ground truth context is provided for each step) for fast inference speed.
AMP implementation.
The implementation is based on the official code in https://github.com/xbpeng/DeepMimic. The encoder consists of three 32-channel convolutional layers interpolated with ReLU activation. The discriminator is implemented as a 3-layer MLP with hidden dimensions of 256 and Tanh activation.
0.A.3 Real-world Offline RL Implementations
We evaluate vision-based offline RL in a real-world Button Press task using a Franka robotic arm with a parallel gripper. We gather 50 expert demonstration videos (rescale to $84\times 84$ resolution) to train reward models for relabeling the offline dataset, which is then used to train policies with the vision-based offline RL method, DrQ+BC [24]. We follow their default hyperparameters. The suboptimal data contains 50% randomly-taken actions.
Appendix 0.B Task Descriptions
We select 7 gripper manipulation tasks from MetaWorld [52] and 3 dexterous manipulation tasks from Adroit [36], as visualized in Figure 12. The tasks are widely used in visual RL and are chosen to be diverse in objects and manipulating skills. All tasks render $64\times 64$-dimensional RGB as the agent’s observation and produce sparse environmental rewards. According to task complexity, we collect 20 expert videos for each MetaWorld task and 50 for each Adroit task. We describe each task below:
@@ Appendix 0.B Task Descriptions
Figure 13: Ablations on diffusion process. The results in (a) suggest that overmuch denoised samples (i.e., 4) may hinder the exploration due to the low variance of estimated entropy. This is further verified in (b) where an appropriate choice of sampling noise scale results in more productive explorations. Results are means of 3 seeds with std. error (shaded area). Red is our default.
Appendix 0.C More Video Prediction Results and Reward Analysis
0.C.1 Video Prediction Results
Qualitative results.
We present the comparison between expert videos and prediction results in Figure 21. We find that the adopted video diffusion model can capture the complex distribution of expert videos from pretraining data and generalize well to unseen expert videos. Interestingly, we also observe that the colored target points in Door Close and Reach are sometimes mispredicted, which may explain the relatively slow exploration at the initial RL training stage, suggesting that a more powerful video diffusion model could be used to improve the reward quality further.
Quantitative analysis.
We compare the video prediction quality between Diffusion Reward and VIPER in terms of three video metrics, SSIM, PSNR, and LPIPS. Results are shown in Table 2, verifying that Diffusion models hold a stronger generalization ability on unseen real robot and simulation trajectories than VideoGPT and thus produce more informative rewards.
0.C.2 Reward Analysis
Figure 14: Reward analysis.
In Figure 3, we conclude that the conclusion that our rewards can generally discern unseen expert behaviors in the MetaWorld domain. We extensively strengthen this conclusion in the Adroit domain and real robot tasks with noisier human-operated demonstrations, as demonstrated in Figure 14. Interestingly, we observe that the suboptimal behaviors are assigned with higher rewards by the diffusion model than by VideoGPT in the Adroit domain, which slightly differs from the observations in the MetaWorld domains. We posit that multiple reasons can be attributed to it. On the one hand, the suboptimal behaviors in the Adroit domain may differ from the ones in MetaWorld due to the higher action space. Besides, the scenes, including background and objects, vary dramatically in the Adorit tasks. Nonetheless, the entropy-based reward performs better than log-likelihood whichever video model is used.
@@ 0.C.2 Reward Analysis
Appendix 0.D More Experiments
0.D.1 More Ablations
We conduct more ablations on our proposed Diffusion Reward in this section. Results are aggregated over Door and Hammer (by default) from Adroit with 3 random seeds.
Sparse environmental reward.
The sparse environmental reward $r^{\mathrm{spar}}$ is integrated into Eq. (5), as the environmental supervision of completion is helpful for RL. We remove $r^{\mathrm{spar}}$ to study its effect. The results in Figure 15 show that the performances decrease dramatically without sparse environmental rewards, indicating that the signal of task completion is necessary for solving complex manipulation tasks. Interestingly, more favorable performance is observed in the Door task. We attribute this to the overlap supervision of RND reward and sparse reward, i.e., exploring novel states (door opening) is partially equivalent to completing the task.
Diffusion process.
Recall that, in Eq. (3), we perform inverse process for $M$ times, and use the generated $M$ samples to estimate the conditional entropy as rewards. The results in Figure 13(a) show that the effect of the number of denoised samples has a slight influence on RL performance when increasing from 1 to 3. However, the learning progress gets stuck in the middle state with 4 denoised samples, though the asymptotic performance is still satisfactory. We posit that this is due to the low variance of estimated entropy, which may lead to more weight on exploitation instead of exploration.
@@ 0.D.1 More Ablations
Meanwhile, we present the time efficiency of different numbers of diffusion processes with a single NVIDIA A40 in Table 3, where the Frames Per Second (FPS) decreases from 87.7 to 45.9 when the number of denoise samples increases from 1 to 4. It suggests that using 1 denoised sample is sufficient to provide informative rewards and retain high inference speed.
@@
Reward coefficient.
We introduce two coefficients to balance the effect of different reward components in Eq. (5). Firstly, we show that using a single coefficient $\alpha$ to balance $\bar{r}^{\mathrm{ce}}$ and $r^{\mathrm{rnd}}$ has no significant distinctions between using two independent ones in Figure 16(a). Secondly, we investigate the effect of the weight of sparse task reward $r^{\mathrm{spar}}$ in Figure 16(b). The results suggest that an appropriate coefficient of task reward $\beta$ is crucial, while a too-low choice may fail to provide sufficient task completion signals to the agent. We set $\beta=1$ throughout the whole paper which is appropriate enough.
Figure 16: Reward coefficient. The results in (a) suggest the introduction of an additional coefficient to balance entropy-based reward and RND reward has a negligible effect but brings more hyperparameter-tuning effort. The results in (b) suggest an appropriate scale of sparse task reward is crucial to give sufficient supervision signals.
Figure 17: Demonstration amount with 3 runs.
Number of demonstrations.
We investigate how sensible our method is to the number of expert videos in Figure 17. Here we collect different numbers of expert videos from three selected Adroit tasks to train our reward models, and test the performance on the most complex Hammer task. Thanks to the strong modeling ability of diffusion models and the appropriate design of entropy-based rewards, the performance is still favorable even given 10 demonstrations for each task.
0.D.2 More Comparison
Figure 18: Comparison with visual representation pretraining on Hammer task with 5 runs.
Comparison with other pretraining schemes.
To further demonstrate the benefit of pre-training our proposed rewards from videos, we do more comparisons with the scheme that pre-trains visual encoders from videos [49, 29, 32]. Specifically, we evaluate two methods, including the representative one that uses PVR [32] to pre-train RL encoder and the one that uses VQ-GAN instead. The results in Figure 18 suggest the remained difficulty of vision-based manipulation tasks even with pretrained visual encoders and, more importantly, the superiority of our reward-pretraining method. Meanwhile, there exist works that directly pre-train decent policies from data for RL [28, 23, 22, 34, 40, 37]. However, they are at the sacrifice of additional collections of reward labels, action labels, or task-agnostic data, while we learn rewards as priors from action-free expert-only videos.
0.D.3 More Generalization Experiments
More tasks.